User Tools

Site Tools


lbaops:lbajun2013:vx017baklog

N.B.: No fringes found and later discovered 100 MHz LO reference not connected! Data not useful.

ASKAP drive problems… potentially… Az drives had servo and brakes red but telescope appeared to perhaps track.

Restarted at 16:37UT. May have lost some time before that.

lbaops/lbajun2013/vx017baklog.txt · Last modified: 2015/12/18 16:38 (external edit)